OpenStudy (anonymous):

The value of.......................? A) 1/3*5+1/5*7+1/7*9+.........+1/23*25 is... Tell me how to solve problem like this..

OpenStudy (asnaseer):

I think this sum can be simplified by using partial fractions, e.g. each term is of the form\[\frac{1}{k(k+2)}=\frac{1}{2}\large(\frac{1}{k}-\frac{1}{k+2}\large)\]so we could get a sum where most terms cancel out

OpenStudy (amistre64):

hmm, i was trying to think how telescopin gwould fit ...

OpenStudy (amistre64):

the thoughts kept going sideways tho

OpenStudy (anonymous):

quite difficult dude

OpenStudy (amistre64):

the telescoping simplifies it nicely

OpenStudy (anonymous):

k

OpenStudy (asnaseer):

@AliBelly have you used partial fractions and telescoping series?

OpenStudy (anonymous):

nope, googling it

OpenStudy (gorv):

\[\frac{ 1 }{ (2k+1)(2k+3) }\]

OpenStudy (gorv):

this will be your series .....now follow the amitsre ' s approch

OpenStudy (gorv):

sorry asnaseer's approch

OpenStudy (asnaseer):

@AliBelly This might be helpful to you: http://www.mathsisfun.com/algebra/partial-fractions.html

OpenStudy (asnaseer):

@gorv expression is the better one to use since all terms have odd numbers. The one I gave only applies for k=3,5,7,... whereas his applies for k=1,2,3,...

OpenStudy (gorv):

yeah i know ..so just for him to take a start tell the equation applicable for his Q

OpenStudy (gorv):

then i said to follow your approch

OpenStudy (anonymous):

\[a _{k}=\frac{ 1 }{ \left( 2k+1 \right)\left( 2k+3 \right) },k=1,2,...,11\] \[\frac{ 1 }{ \left( 2k+1 \right) \left( 2k+3 \right)}=\frac{ 1 }{ \left( 2k+1 \right)\left( -1+3 \right)}+\frac{ 1 }{ \left( -3+1 \right)\left( 2k+3 \right) }\] \[=\frac{ 1 }{ 2 }\left[ \frac{ 1 }{ 2k+1 }-\frac{ 1 }{ 2k+3 } \right]\] put k=1,2,...,11 and add vertically downwards